OpenStudy (anonymous):

Find the distance EA for the points E(–8, –2) and A(4, 1).

OpenStudy (anonymous):

The answer is \[\sqrt{(-8-4)^{2} + (-2-1)^{2}}\] Which is 12.37

OpenStudy (anonymous):

What are the dimensions of a rectangle that has a perimeter of 74 units and an area of 102 square units?

OpenStudy (anonymous):

2x + 2y = 74 xy = 102 2(102/y) + 2y = 74 Solve for y and plug that into an equation to solve for x(:

OpenStudy (anonymous):

Express each ratio as a unit rate. Round to the nearest tenth, if necessary. $3 for 6 cans of tuna